# A Survey on Knowledge Organization Systems of Research Fields: Resources and Challenges

### v1

@@ 4.2.3 Depth
Similarly to the number of concepts, the depth of KOSs also displays considerable variety, ranging from 1 (Web of Science Categories) to 39 (Open Biological and Biomedical Ontology). In particular, 15 KOSs employ up to 3 levels, 15 use from 4 to 6 levels, 3 utilise from 7 to 10 levels, and 11 extend beyond 10 levels.
@@ 4.2.4 Kind of hierarchy
Column ‘Hr’ in Table 1 reveals a near-equal distribution between poly-hierarchical (24) and mono-hierarchical (20) KOSs. Only one KOS (Web of Science Categories) is considered non-hierarchical since it is a flat list of terms. When examining the 23 single-field KOSs, it is interesting to note that poly-hierarchical structures (14) outnumber mono-hierarchical ones (9). In contrast, multi-field systems exhibit a nearly equal distribution, with 11 mono-hierarchical and 10 poly-hierarchical structures. This trend suggests that single-field KOSs are more likely to employ poly-hierarchical structures, as they need to capture fine-grained research topics that may be derived from multiple parent topics.
@@ 4.2.5 Related terms
Noteworthy exceptions are the Mathematics Subject Classification and the Dewey Decimal Classification. Although the Mathematics Subject Classification is traditionally considered a taxonomy (Dunne and Hulek, 2020), it also includes related terms. This functionality is enabled on the zbMATH website181818zbMATH - https://zbmath.org/classification/, where the KOS is displayed with a “see also” anchor, without adherence to the ANSI/NISO standard (ANSI/NISO Z39.19-2005, R2010). Similarly, the Dewey Decimal Classification uses the “see-also” references for synonyms and related terms (Mitchell, 2001).
@@ 4.3.1 Generation
Only two KOSs are generated using a fully automated pipeline, both within the field of computer science: the Computer Science Ontology and Computer Science Subject Headings from Wikipedia. The Computer Science Ontology was generated using Klink-2 (Osborne and Motta, 2015), which processed 16 million scientific publications. Klink-2 identifies relationships between topics by analysing various indicators, including co-occurrence patterns, temporal distributions, and label similarity. The Computer Science Subject Headings from Wikipedia was created through an automated approach aimed at enhancing and refining the “Computer Science” branch of the Wikipedia category system. This approach integrates community detection, machine learning, and manually developed heuristics to identify and incorporate additional topics from Wikipedia articles Han et al. (2020).
@@ 4.3.2 Distribution formats
The majority of KOSs (27) are accessible in HTML format, primarily through their providers’ websites. In this format, the KOS is essentially a webpage listing and connecting concepts. While this is human-readable, the underlying data lacks the structured organisation necessary for direct and automated processing by computer systems. Several KOSs are available in machine-readable formats: 17 utilise the RDF standard, and 10 are available in CSV format. Finally, 6 KOSs are exclusively available in PDF format, which complicates their integration with other KOSs and hinders automatic analysis.
@@ 4.3.4 Frequency of updates
We also investigated whether the frequency of updates is related to the type of system. Our analysis revealed that among the 18 KOSs updated within a year, the majority were ontologies (11), followed by taxonomies (5) and thesauri (2). In contrast, the 8 KOSs that were either updated every ten years or discontinued were taxonomies. This suggests that while taxonomies might be effective initially for organising subjects hierarchically, they may become increasingly challenging to maintain over time due to the growing complexity and unique characteristics of the subjects they represent. On the other hand, ontologies, often implemented in flexible formats such as OWL and RDF, may be easier to maintain.
@@ 4.3.5 Last Update
Seventeen systems have received their most recent updates in 2023 or later, while thirteen were last updated between 2020 and 2022. This indicates that within the last five years, thirty KOSs have been revised to reflect the evolving nature of their respective fields. Of these, 14 are single-field systems, and 15 are multi-field.
The remaining fifteen systems either received their last update before 2020 or their latest release date could not be determined. Notably, key disciplines such as “Chemistry” (with ChemOnt, last updated in 2016) and “Geology” (with the U.S. Geological Survey Library Classification System, last updated in 2000, and the GeoRef Thesaurus, last updated in 2008) can only rely on possibly outdated KOSs. This is potentially a significant issue, as the outdated representations in these disciplines can hinder the dissemination of modern research efforts that address topics not fully covered by these KOSs.
@@ 4.3.6 Languages
Our exclusion criteria ensured that all KOSs had at least one version in English. Thirty-five KOSs are available exclusively in English. Conversely, ten KOSs provide research topics in multiple languages, driven by their intended use and jurisdictional context. For example, EuroVoc, developed by the Publications Office of the European Union, supports 23 languages to facilitate seamless interoperability within European digital libraries. The additional nine multilingual KOSs are Agrovoc Thesaurus, Dewey Decimal Classification, Fields of research and development, STW Thesaurus for Economics, Unified Medical Language System, TheSoz, Science Metrix Classification, Unesco Thesaurus and Art and Architecture Thesaurus.
@@
Furthermore, KOSs often vary in the number of concepts available in different languages. Typically, non-English versions only provide a partial representation of the domain described by the English version. For instance, as shown in Fig. 3, Agrovoc provides good coverage in English, French, Turkish, Spanish, Arabic, and a few other languages. However, the number of available concepts is significantly limited in languages such as Estonian, Burmese, Khmer, and Greek. Naturally, achieving uniform coverage poses greater challenges for KOSs with both a high volume of concepts and extensive multilingual support. A few smaller KOSs, such as Eurovoc, Unesco Thesaurus, Fields of Research and Development, Science Metrix Classification, and STW Thesaurus for Economics, have managed to maintain consistent representation across all supported languages.
@@ 4.3.7 Providers and Maintainers
National Libraries, including the US National Library of Medicine, the Library of Congress, the German National Library of Economics (ZBW), and the German National Library of Science and Technology (TIB) provided 7 KOS.
@@ 4.4 External Links
Eighteen KOSs provide links to external knowledge bases. This is typically done by defining mappings that indicate that two entities in different representations refer to the same real-world object or concept. They typically rely on well-know semantic relationships such as owl:same_as, skos:exactMatch, and skos:closeMatch. As an example, the concept “sunflowers” available in Agrovoc Thesarus [^21] Sunflowers (plural) in Agrovoc Thesarus - http://aims.fao.org/aos/agrovoc/c˙aad037e4 has a skos:closeMatch with “sunflower” in Eurovoc [^22] Sunflower (singular) in Eurovoc - http://eurovoc.europa.eu/4472.
The KOSs under analysis typically connect either to general knowledge graphs, such as Wikidata or DBpedia, or to other KOSs. Such interconnections are generally beneficial, as they create a network of related resources, providing diverse perspectives on research topics and facilitating the creation of novel resources.
Wikidata [^23] Wikidata - https://www.wikidata.org stands out as the most externally linked knowledge graph since it is reached by concepts from Agrovoc Thesaurus, Computer Science Ontology, STW Thesaurus for Economics, EuroVoc, Open Biological and Biomedical Ontology, and the Library of Congress Class. (and Subj. Head.). Conversely, Agrovoc Thesaurus, ChemOnt, STW Thesaurus for Economics, EuroVoc are the KOSs with the highest number of externally connected knowledge bases. For instance, the AGROVOC Thesaurus is highly interconnected, linking its concepts to EuroVoc, Wikidata, DBpedia, the UNESCO Thesaurus, and the Library of Congress Subject Headings (LCSH). Similarly, EuroVoc is linked not only with the AGROVOC Thesaurus, but also with TheSoz, LCSH, UNESCO, STW, MeSH, Wikidata, and several other knowledge organization systems.
Twelve out of the initial eighteen earn a five-star rating within the Linked Open Data deployment scheme proposed by Sir Tim Berners-Lee in 2010242424Linked Data - https://www.w3.org/DesignIssues/LinkedData.html. Their extensive interconnections, facilitated by RDF standards such as OWL and SKOS, establish a rich network of information with related KOSs. Adhering to the 5-star scheme [^25] Linked Open Data 5 Star - https://www.w3.org/DesignIssues/LinkedData.html optimises discoverability, reusability, and the potential for integration into the broader knowledge web, fostering innovative applications and the development of more comprehensive knowledge organization systems.
@@ 4.5 Usage
Various KOS are also utilised across a wide array of initiatives to directly advance research and create additional knowledge and tools. For instance, the Computer Science Ontology has been employed for exploring and analysing scholarly data through platforms like Rexplore (Osborne et al., 2013), ScholarLensViz (Löffler et al., 2020), ConceptScope (Zhang et al., 2021), and VeTo (Vergoulis et al., 2020). It has also been instrumental in generating knowledge graphs, such as AIDA KG (Angioni et al., 2022b) and CS KG (Dessì et al., 2022), as well as in recommending video lessons (Borges and dos Reis, 2019). However, thoroughly analysing these specific applications would require a dedicated survey, as exemplified by Jing’s usage analysis of the Unified Medical Language System (Jing, 2021).
@@ 5.1 Towards a comprehensive and granular representation of all academic disciplines
We argue that such a resource would be transformative in this field and allow: i) organising content across various digital libraries, thereby enhancing interoperability; ii) facilitating the retrieval and analysis of research outputs (e.g., articles, patents, project reports) and agents (e.g., researchers, organisations); iii) enabling the monitoring of research, development, and innovation activities; iv) ensuring data quality and compatibility across research institutions; and v) supporting evidence-based policymaking.
A potential approach to creating this resource is to adopt one or more multi-field KOSs to represent broad research areas, and then integrate several single-field KOSs to effectively cover specific disciplines. As a first step, this section will explore which KOSs could be integrated to develop a comprehensive system. In Section 5.2, we will discuss current methods for interlinking multiple KOSs.
@@
It is useful to analyse to what extent the current multi-field KOSs fulfil the requirements defined in the previous session. Table 4 presents an evaluation of the 22 multi-field KOSs examined in this study, considering four key features: comprehensiveness, depth, frequency of updates, and openness. For comprehensiveness, colours are assigned based on the KOS coverage, detailed by Table 2. Specifically, green indicates full coverage, orange denotes partial coverage, and red signifies that the KOSs cover only a limited set of fields. In the depth column, the colour coding is as follows: red indicates a depth of 2 or less (low or non-existent), orange represents a depth between 3 and 5 (medium), and green signifies a depth of 6 or higher (high). In the frequency of update column, red indicates that the KOS is discontinued with no future updates planned, orange is used for slow updates (i.e., every 5–10 years), and green denotes more frequent updates (i.e., within 2 years). If such information is unavailable, it is marked in white. Lastly, in the openness column, red indicates that the KOS is copyrighted and requires a license, orange signifies that the KOS can be used with certain restrictions, and green shows that it is an open resource.
Table 4: Analysis of KOSs according to their comprehensiveness, depth, frequency of update, and openness. Features of each KOSs are evaluated using a colour-coded system: $\blacksquare$ green indicates a strong performance, $\blacksquare$ orange signifies an adequate performance, and $\blacksquare$ red reflects poor performance. A feature is blank when the information is not available.
@@
Only five KOSs include all 19 broad fields of study: All Science Journal Classification Codes, Fields of Research (ANZSRC), Research Fields, Courses and Disciplines (ASRC), Micorsoft’s Fields of Study, and Web of Science Categories. However, none of these KOSs fulfils all the other requirements. All Science Journal Classification Codes and Web of Science Categories are very shallow and not open. The Fields of Research, Research Fields, Courses and Disciplines (ASRC), and the Microsoft’s Fields of Study offer more depth compared to the previous two, but they were discontinued. The Fields of Research (ANZSRC) is limited in depth, with many specific categories still overly broad, and updates are infrequent, often taking several years to implement. The Dewey Decimal Classification and Library of Congress Subject Headings offer extensive coverage across nearly all areas, along with substantial depth and frequent updates. However, their openness remains an issue, as users need to pay subscription fees for using them.
@@
Several disciplines lack dedicated KOS and are only superficially addressed by multi-field systems, making it challenging to categorise documents with the necessary precision. In particular, we were unable to identify any KOS that offers a good characterisation of seven research fields: “History”, “Political Science”, “Environmental Science”, “Material Science”, “Geography”, “Sociology”, and “Business”. In the fields of “Political Science” and “Sociology”, The Soz could potentially be utilised. However, The Soz mostly focuses on “Social Science” and does not fully cover these two disciplines. The absence of fine-grained representation for these seven major areas is notable and underscores the need for further development in this space. Automated methods for generating KOSs can offer a valuable solution in this regard, a topic we will further discuss in Section 5.7.
@@ 5.2 Integration of multiple KOSs
In the following, we discuss future directions regarding the integration of multiple KOSs, which may lead to the creation of a more comprehensive and granular representation of research fields.
@@ 3. Sub-challenge 3:
Automatic approaches typically use a combination of similarity metrics, natural language processing, and machine learning (Salatino et al., 2020; Zapilko et al., 2013; Declerck, 2013). For instance, Salatino et al. (2020) associated research topics to the corresponding DBpedia entity with the DBpedia Spotlight API (Daiber et al., 2013). They fed the tool with artificial sentences listing the labels of the topic and of its direct sub- and super-topics, and then it returned the related DBpedia entities alongside the similarity score. Zapilko et al. (2013) mapped TheSoz to Agrovoc and DBpedia, using string similarity between terms and retaining the matches that have Levenshtein distance lower than a threshold. In the domain of KOSs of academic fields, we still typically rely on simple approaches, which mainly use lexical heuristics, and may lead to three potentially unintended consequences (Shvaiko and Euzenat, 2008; Solimando et al., 2014; Slater et al., 2020). First, the mapping might introduce new internal relationships between the entities of one system, and therefore modify inadvertently the description of the domain. Instead, the mapping should just enable the interaction across KOSs. Second, the automatic integration can introduce logical inconsistencies. Finally, the mapping might connect entities belonging to different contexts, indicating a potential mapping error (Jiménez-Ruiz et al., 2011). Some recent approaches address these limitations by relying on description logic (Dhombres and Bodenreider, 2016) or deep learning (Yip et al., 2019). For instance, Dhombres and Bodenreider (2016) developed an approach for mapping Human Phenotype Ontology (HPO) and the Standardized Nomenclature of Medicine Clinical Terms (SNOMED CT) using both lexical and logical approaches. The latter approach consists of using a representation based on description logic for comparing the concepts. This method can mitigate the limitations of the lexical approach, however not all KOSs are developed through description logic. Yip et al. (2019) developed a deep learning model with a Siamese recurrent architecture to identify synonyms across UMLS concepts. The model provides good results; however, it still requires more research and fine-tuning because it presents several false positives (matches together non-synonyms) and false negatives (fails to identify synonyms). The community still needs to further investigate and develop these new solutions in practical settings.
@@
Based on the 5-star deployment scheme for open data2929295-star Open Data — https://www.w3.org/DesignIssues/LinkedData.html, among the possible publishing formats, RDF is the one that fosters better integration. A system published according to this standard consists of a set of RDF statements. Each statement is a three-part structure (also known as triple) which is the smallest irreducible representation for binary relationships, and it is expressed in the form of <subject, predicate, object(Berners-Lee et al., 2001). For instance, <Social Sciences, narrower, Politics> indicates that Social Sciences is a broader area of Politics. In this way, two entities (both subject and object) are linked via a predicate or verb. In addition, every part of a triple is individually addressable through unique URIs, such as: <http://www.nature.com/subjects/social-sciences, http://www.w3.org/2004/02/skos/core#narrower, http://www.nature.com/subjects/politics>. Such representation allows AI systems to interconnect, identify, disambiguate, and integrate data effectively.
On the contrary, KOSs published in CSV or HTML are harder to treat since the relevant knowledge is not described in a structured format, and relations may be implicit or interpretable only in a wider context. Generating a representation of these KOSs in a standard format such as RDF is not a trivial task. For example, when working with CSV files, it is essential to understand the structure and data types of the columns (e.g., string, integer, float, date). Tools such as RML (Dimou et al., 2014) can be used to automate the conversion process from CSV to RDF. However, this process still requires careful consideration of the data’s format and structure to ensure accurate transformation. Similarly, KOSs published in HTML often have arbitrary structures, making it necessary to develop custom parsers to convert them into RDF. When KOSs are available only as PDF files, parsing and extracting the information becomes even more challenging.
In conclusion, the RDF format is adopted by only half (11 out of 23) of the single-field KOSs across the different disciplines, as shown in Table 3. Several significant academic fields (e.g.,“Psychology”, “Chemistry”, “Geology”, “Engineering”, “Philosophy”) do not yet rely on standard machine-readable formats, hindering the reuse and integration of their KOSs. In the literature, we can find a few research efforts to RDFy [^30] RDFication refers to the process of converting data into Resource Description Framework (RDF) format. KOSs. Examples include the RDFification of the Mathematics Subject Classification [^31] MSC2020_SKOS (TIBHannover) - https://github.com/TIBHannover/MSC2020_SKOS and IEEE Thesaurus323232ieee-taxonomy-thesaurus-rdf (The Open University) - https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf. However, these RDF versions are not maintained by the original curators of the KOSs, making them difficult to update and reuse.
@@
The first KOSs were of relatively small size, such as the Library Classification for Environmental Science [^33] We excluded the Library Classification for Environmental Science from our analysis because it was developed over five decades ago and is no longer actively maintained or used by the scientific community., developed by Plate (1966). Typically, their creation would involve sketching ideas on paper or writing down topics on sticky notes, which would then be arranged on a table to create a hierarchical structure (Motta, 1999). Today, we have access to a wide range of tools that streamline this process, enabling the creation of more structured and complex KOSs. For instance, the German Centre for Higher Education Research and Science Studies [^34] German Centre for Higher Education Research and Science Studies (DZHW) — www.dzhw.eu (DZHW) used Trello to build the Research Core Dataset [^35] Research Core Dataset — https://w3id.org/kdsf-ffk. Despite being a KOS of academic disciplines, RCD was not included in our analysis in Sec. 4 because it is available only in German (see selection criteria in Sec. 3.2). (RCD), a classification of interdisciplinary research fields (Stiller et al., 2021). Trello [^36] Trello — https://trello.com is a web-based project management tool implementing the Kanban system, and it allows multiple users to collaborate on the same board. To build RCD, the experts created a card for each subject and then arranged them over the board to create the hierarchical structure. However, building complex KOSs using Trello presents some limitations, since it does not allow users to nest cards in more than two levels. In addition, as Trello is not intended for this task, it does not provide any tool to handle disagreement among the experts due to their different backgrounds.
@@
More advanced tools for building KOSs are Protégé [^37] Protégé — https://protege.stanford.edu, Semantic MediaWiki [^38] Semantic MediaWiki — https://www.semantic-mediawiki.org and PoolParty [^39] PoolParty — https://www.poolparty.biz. Protégé is an open-source software developed by the Stanford University to support developers in creating reusable ontologies and building knowledge-based systems (Musen, 2015). Its graphical interface offers a range of functionalities for browsing and editing ontologies. While the original Protégé software lacks native collaborative features, its cloud-based counterpart, WebProtégé, enables multiple users to work simultaneously on ontology development (Tudorache et al., 2013).

### v2

@@ 4.2.3 Depth
Similarly to the number of concepts, the depth of KOSs also displays considerable variety, ranging from 1 (Web of Science Categories) to 39 (Open Biological and Biomedical Ontology).
@@ 4.2.4 Kind of hierarchy
Column ‘Hr’ in Table 1 reveals a near-equal distribution between poly-hierarchical (24) and mono-hierarchical (20) KOSs. Only one KOS (Web of Science Categories) is considered non-hierarchical since it is a flat list of terms.
@@ 4.2.5 Related terms
Noteworthy exceptions are the Mathematics Subject Classification and the Dewey Decimal Classification. Although the Mathematics Subject Classification is traditionally considered a taxonomy (Dunne and Hulek, 2020), it also includes related terms. This functionality is enabled on the zbMATH website181818zbMATH — https://zbmath.org/classification, where the KOS is displayed with a “see also” anchor, without adherence to the ANSI/NISO standard (ANSI/NISO Z39.19-2005, R2010). Similarly, the Dewey Decimal Classification uses the “see-also” references for synonyms and related terms (Mitchell, 2001).
@@ 4.3.1 Generation
Only two KOSs are generated using a fully automated pipeline, both within the field of computer science: the Computer Science Ontology and Computer Science Subject Headings from Wikipedia. The Computer Science Ontology was generated using Klink-2 (Osborne and Motta, 2015), which processed 16 million scientific publications. Klink-2 identifies relationships between topics by analysing various indicators, including co-occurrence patterns, temporal distributions, and label similarity. The Computer Science Subject Headings from Wikipedia was created through an automated approach aimed at enhancing and refining the “Computer Science” branch of the Wikipedia category system. This approach integrates community detection, machine learning, and manually developed heuristics to identify and incorporate additional topics from Wikipedia articles (Han et al., 2020).
@@ 4.3.2 Distribution formats
The majority of KOSs (27) are accessible in HTML format, primarily through their providers’ websites. In this format, the KOS is essentially a webpage listing and connecting concepts. While this is human-readable, the underlying data lacks the structured organisation necessary for direct and automated processing by computer systems. Several KOSs are available in machine-interpretable formats: 17 utilise the RDF standard, and 10 are available in CSV format. Finally, 6 KOSs are exclusively available in PDF format, which complicates their integration with other KOSs and hinders automatic analysis.
@@ 4.3.4 Frequency of updates
Figure 4:
@@ 4.3.5 Last Update
Twenty systems have received their most recent update in 2023 or later, while thirteen were last updated between 2020 and 2022. This indicates that within the last five years, thirty-three KOSs have been revised to reflect the evolving nature of their respective fields. Of these, 17 are single-field systems, and 15 are multi-field. The remaining twelve systems either received their last update before 2020 or their latest release date could not be determined.
Notably, key disciplines such as “Chemistry” (with ChemOnt, last updated in 2016) and “Geology” (with the U.S. Geological Survey Library Classification System, last updated in 2000, and the GeoRef Thesaurus, last updated in 2008) can only rely on possibly outdated KOSs. This is potentially a significant issue, as the outdated representations in these disciplines can hinder the dissemination of modern research efforts that address topics not fully covered by these KOSs.
@@ 4.3.6 Languages
Furthermore, KOSs often vary in the number of concepts available in different languages. Typically, non-English versions only provide a partial representation of the domain described by the English version. For instance, as shown in Figure 5, Agrovoc provides good coverage in English, French, Turkish, Spanish, Arabic, and a few other languages. However, the number of available concepts is significantly limited in languages such as Estonian, Burmese, Khmer, and Greek. Naturally, achieving uniform coverage poses greater challenges for KOSs with both a high volume of concepts and extensive multilingual support.
@@ 4.3.7 Providers and Maintainers
National Libraries, including the US National Library of Medicine, the Library of Congress, the German National Library of Economics (ZBW), and the German National Library of Science and Technology (TIB) provided 7 KOSs.
@@ 4.4 External Links
Nineteen KOSs provide links to external knowledge bases. This is typically done by defining mappings that indicate that two entities in different representations refer to the same real-world object or concept. They typically rely on well-know semantic relationships such as owl:same_as, skos:exactMatch, and skos:closeMatch. As an example, the concept “sunflowers” available in Agrovoc Thesarus [^21] Sunflowers (plural) in Agrovoc Thesarus — http://aims.fao.org/aos/agrovoc/c˙aad037e4 has a skos:closeMatch with “sunflower” in Eurovoc [^22] Sunflower (singular) in Eurovoc — http://eurovoc.europa.eu/4472.
The KOSs under analysis typically connect either to general knowledge graphs (Peng et al., 2023), such as Wikidata [^23] Wikidata — https://www.wikidata.org or DBpedia [^24] DBpedia — https://www.dbpedia.org/, or to other KOSs. Such interconnections are generally beneficial, as they create a network of related resources, providing diverse perspectives on research topics and facilitating the creation of novel resources.
Wikidata stands out as the most externally linked knowledge graph since it is reached by concepts from Agrovoc Thesaurus, Computer Science Ontology, STW Thesaurus for Economics, EuroVoc, Open Biological and Biomedical Ontology, and the Library of Congress Class. (and Subj. Head.). Conversely, Agrovoc Thesaurus, ChemOnt, STW Thesaurus for Economics, EuroVoc are the KOSs with the highest number of externally connected knowledge bases. For instance, the AGROVOC Thesaurus is highly interconnected, linking its concepts to EuroVoc, Wikidata, DBpedia, the UNESCO Thesaurus, and the Library of Congress Subject Headings (LCSH). Similarly, EuroVoc is linked not only with the AGROVOC Thesaurus, but also with TheSoz, LCSH, UNESCO, STW, MeSH, Wikidata, and several other knowledge organization systems.
Thirteen out of the initial nineteen earn a five-star rating within the Linked Open Data deployment scheme proposed by Sir Tim Berners-Lee in 2010252525Linked Data — https://www.w3.org/DesignIssues/LinkedData.html. Their extensive interconnections, facilitated by RDF standards such as OWL and SKOS, establish a rich network of information with related KOSs. Adhering to the 5-star scheme [^26] Linked Open Data 5 Star — https://www.w3.org/DesignIssues/LinkedData.html optimises discoverability, reusability, and the potential for integration into the broader knowledge web, fostering innovative applications and the development of more comprehensive knowledge organization systems.
@@ 4.5 Usage
Various KOSs are also utilised across a wide array of initiatives to directly advance research and create additional knowledge and tools. For instance, the Computer Science Ontology has been employed for exploring and analysing scholarly data through platforms like Rexplore (Osborne et al., 2013), ScholarLensViz (Löffler et al., 2020), ConceptScope (Zhang et al., 2021), and VeTo (Vergoulis et al., 2020). It has also been instrumental in generating knowledge graphs, such as AIDA KG (Angioni et al., 2022b) and CS KG (Dessì et al., 2022), as well as in recommending video lessons (Borges and dos Reis, 2019). However, thoroughly analysing these specific applications would require a dedicated survey, as exemplified by Jing’s usage analysis of the Unified Medical Language System (Jing, 2021).
@@ 5.1 Towards a comprehensive and granular representation of all academic disciplines
We argue that such a resource would be transformative in this field and allow: i) organising content across various digital libraries, thereby enhancing interoperability; ii) facilitating the retrieval and analysis of research outputs (e.g., articles, patents, project reports) and agents (e.g., researchers, organisations); iii) enabling the monitoring of research, development, and innovation activities; iv) ensuring data quality and compatibility across research institutions; and v) supporting evidence-informed policymaking.
A potential approach to creating this resource is to adopt one or more multi-field KOSs to represent broad research areas and then integrate several single-field KOSs to effectively cover specific disciplines. As a first step, this section will explore which KOSs could be integrated to develop a comprehensive system. In Section 5.2, we will discuss current methods for interlinking multiple KOSs.
@@
It is useful to analyse to what extent the current multi-field KOSs fulfil the requirements defined in the previous session. Table 6 presents an evaluation of the 22 multi-field KOSs examined in this study, considering four key features: comprehensiveness, depth, frequency of updates, and openness.
Table 6:
@@
Only five KOSs include all 19 broad fields of study: All Science Journal Classification Codes, Fields of Research (ANZSRC), Research Fields, Courses and Disciplines (ASRC), Micorsoft’s Fields of Study, and Web of Science Categories. However, none of these KOSs fulfils all the other requirements. All Science Journal Classification Codes and Web of Science Categories are very shallow and not open. The Fields of Research, Research Fields, Courses and Disciplines (ASRC), and the Microsoft’s Fields of Study offer more depth compared to the previous two, but they were discontinued. The Fields of Research (ANZSRC) is limited in depth, with many specific categories still overly broad, and updates are infrequent, often taking several years to implement. The Dewey Decimal Classification and Library of Congress Subject Headings offer extensive coverage across nearly all areas, along with substantial depth and frequent updates. However, their openness remains an issue, as users need to pay subscription fees to use them.
@@
Several disciplines lack dedicated KOSs and are only superficially addressed by multi-field systems, making it challenging to categorise documents with the necessary precision. In particular, we were unable to identify any KOS that offers a good characterisation of seven research fields: “History”, “Political Science”, “Environmental Science”, “Material Science”, “Geography”, “Sociology”, and “Business”. In the fields of “Political Science” and “Sociology”, The Soz could potentially be utilised. However, The Soz mostly focuses on “Social Science” and does not fully cover these two disciplines. The absence of fine-grained representation for these seven major areas is notable and underscores the need for further development in this space. Automated methods for generating KOSs can offer a valuable solution in this regard, a topic we will further discuss in Section 5.7.
@@ 5.2 Integration of multiple KOSs
In the following, we discuss future directions regarding the integration of multiple KOSs, which may lead to creating a more comprehensive and granular representation of research fields.
@@ 3. Sub-challenge 3:
Automatic approaches typically use a combination of similarity metrics, natural language processing, and machine learning (Salatino et al., 2020b; Zapilko et al., 2013; Declerck, 2013). For instance, Salatino et al. (2020b) associated research topics to the corresponding DBpedia entity with the DBpedia Spotlight API (Daiber et al., 2013). They fed the tool with artificial sentences listing the labels of the topic and of its direct sub- and super-topics, and then it returned the related DBpedia entities alongside the similarity score. Zapilko et al. (2013) mapped TheSoz to Agrovoc and DBpedia, using string similarity between terms and retaining the matches that have Levenshtein distance lower than a threshold. In the domain of KOSs of academic fields, we still typically rely on simple approaches, which mainly use lexical heuristics, and may lead to three potentially unintended consequences (Shvaiko and Euzenat, 2008; Solimando et al., 2014; Slater et al., 2020). First, the mapping might introduce new internal relationships between the entities of one system, and therefore modify inadvertently the description of the domain. Instead, the mapping should just enable the interaction across KOSs. Second, the automatic integration can introduce logical inconsistencies. Finally, the mapping might connect entities belonging to different contexts, indicating a potential mapping error (Jiménez-Ruiz et al., 2011). Some recent approaches address these limitations by relying on description logic (Dhombres and Bodenreider, 2016) or deep learning (Yip et al., 2019). For instance, Dhombres and Bodenreider (2016) developed an approach for mapping Human Phenotype Ontology (HPO) and the Standardized Nomenclature of Medicine Clinical Terms (SNOMED CT) using both lexical and logical approaches. The latter approach consists of using a representation based on description logic to compare the concepts. This method can mitigate the limitations of the lexical approach, however, not all KOSs are developed through description logic. Yip et al. (2019) developed a deep learning model with a Siamese recurrent architecture to identify synonyms across UMLS concepts. The model provides good results; however, it still requires more research and fine-tuning because it presents several false positives (matches together non-synonyms) and false negatives (fails to identify synonyms). The community still needs to further investigate and develop these new solutions in practical settings.
@@
Based on the 5-star deployment scheme for open data3030305-star Open Data — https://www.w3.org/DesignIssues/LinkedData.html, among the possible publishing formats, RDF is the one that fosters better integration. A system published according to this standard consists of a set of RDF statements. Each statement is a three-part structure (also known as triple) which is the smallest irreducible representation for binary relationships, and it is expressed in the form of <subject, predicate, object(Berners-Lee et al., 2001). For instance, <Social Sciences, narrower, Politics> indicates that Social Sciences is a broader area of Politics. In this way, two entities (both subject and object) are linked via a predicate or verb. In addition, every part of a triple is individually addressable through unique URIs, such as: <http://www.nature.com/subjects/social-sciences, http://www.w3.org/2004/02/skos/core#narrower, http://www.nature.com/subjects/politics>. Such representation allows AI systems to interconnect, identify, disambiguate, and integrate data effectively.
On the contrary, KOSs published in CSV or HTML are more challenging to treat since the relevant knowledge is not described in a structured format, and relations may be implicit or interpretable only in a wider context. Generating a representation of these KOSs in a standard format such as RDF is not a trivial task. For example, when working with CSV files, it is essential to understand the structure and data types of the columns (e.g., string, integer, float, date). Tools such as RML (Dimou et al., 2014) can be used to automate the conversion process from CSV to RDF. However, this process still requires careful consideration of the data’s format and structure to ensure accurate transformation. Similarly, KOSs published in HTML often have arbitrary structures, making it necessary to develop custom parsers to convert them into RDF. When KOSs are available only as PDF files, parsing and extracting the information becomes even more challenging.
In conclusion, the RDF format is adopted by only half (11 out of 23) of the single-field KOSs across the different disciplines, as shown in Table 4. Several significant academic fields (e.g.,“Psychology”, “Chemistry”, “Geology”, “Engineering”, “Philosophy”) do not yet rely on standard machine-readable formats, hindering the reuse and integration of their KOSs. In the literature, we can find a few research efforts to RDFy [^31] RDFication refers to the process of converting data into Resource Description Framework (RDF) format. KOSs. Examples include the RDFification of the Mathematics Subject Classification [^32] MSC2020_SKOS (TIBHannover) — https://github.com/TIBHannover/MSC2020_SKOS and IEEE Thesaurus333333ieee-taxonomy-thesaurus-rdf (The Open University) — https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf. However, these RDF versions are not maintained by the original curators of the KOSs, making them difficult to update and reuse.
@@
The first KOSs were of relatively small size, such as the Library Classification for Environmental Science [^34] We excluded the Library Classification for Environmental Science from our analysis because it was developed over five decades ago and is no longer actively maintained or used by the scientific community., developed by Plate (1966). Typically, their creation would involve sketching ideas on paper or writing down topics on sticky notes, which would then be arranged on a table to create a hierarchical structure (Motta, 1999). Today, we have access to a wide range of tools that streamline this process, enabling the creation of more structured and complex KOSs. For instance, the German Centre for Higher Education Research and Science Studies [^35] German Centre for Higher Education Research and Science Studies (DZHW) — www.dzhw.eu (DZHW) used Trello to build the Research Core Dataset [^36] Research Core Dataset — https://w3id.org/kdsf-ffk. Despite being a KOS of academic disciplines, RCD was not included in our analysis in Sec. 4 because it is available only in German (see selection criteria in Sec. 3.2). (RCD), a classification of interdisciplinary research fields (Stiller et al., 2021). Trello [^37] Trello — https://trello.com is a web-based project management tool implementing the Kanban system, and it allows multiple users to collaborate on the same board. To build RCD, the experts created a card for each subject and then arranged them over the board to create the hierarchical structure. However, building complex KOSs using Trello presents some limitations, since it does not allow users to nest cards on more than two levels. In addition, as Trello is not intended for this task, it does not provide any tool to handle disagreement among the experts due to their different backgrounds.
@@
More advanced tools for building KOSs are Protégé [^38] Protégé — https://protege.stanford.edu, Semantic MediaWiki [^39] Semantic MediaWiki — https://www.semantic-mediawiki.org, VocBench3404040VocBench — https://vocbench.uniroma2.it, and PoolParty [^41] PoolParty — https://www.poolparty.biz. Protégé is an open-source software developed by Stanford University to support developers in creating reusable ontologies and building knowledge-based systems (Musen, 2015). Its graphical interface offers a range of functionalities for browsing and editing ontologies. While the original Protégé software lacks native collaborative features, its cloud-based counterpart, WebProtégé, enables multiple users to work simultaneously on ontology development (Tudorache et al., 2013).